OpenStudy (anonymous):

Bruce pitches for his little brother's baseball team. He has observed that the number of pitches a batter hits varies and is given by the function f(x) = x – 11, where x is the number of pitches and f(x) is the number of hits. If batters get {4, 12, 14, 27, 42} hits, how many pitches did Bruce throw? {13, 21, 24, 26, 44} {17, 24, 27, 28, 35} {20, 27, 32, 33, 45} {15, 23, 25, 38, 53}

OpenStudy (michele_laino):

for example I conside 4 hits, so I can write this equation: 4 = x -11

OpenStudy (michele_laino):

consider*

OpenStudy (michele_laino):

now I add 11 to both sides, so i get: 4+11=x-11 +11 please simplify and solve for x

OpenStudy (anonymous):

x=15

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

ok so it's D

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (anonymous):

ok thanks

OpenStudy (michele_laino):

:)